OpenStudy (anonymous):

If a cone and a hemisphere have the same radius and the same volume, what is the height of the cone in terms of the radius?

OpenStudy (perl):

Volume of a cone = 1/3 * Pi * r^2 * h volume of hemisphere = 1/2 * (4/3 * pi * r^3 )

OpenStudy (perl):

so we can set them equal to each other and solve for h in terms of r